# Multi-Continental Healthcare Modelling Using Blockchain-Enabled Federated Learning

### v2

@@ Abstract
One of the biggest challenges of building artificial intelligence (AI) model in healthcare area is the data sharing. Since healthcare data is private, sensitive, and heterogeneous, collecting sufficient data for modelling is exhausted, costly, and sometimes impossible. In this paper, we propose a framework for global healthcare modelling using datasets from multi-continents (Europe, North America and Asia) while without sharing the local datasets, and choose glucose management as a study model to verify its effectiveness. Technically, blockchain-enabled federated learning is implemented with adaption to make it meet with the privacy and safety requirements of healthcare data, meanwhile rewards honest participation and penalize malicious activities using its on-chain incentive mechanism. Experimental results show that the proposed framework is effective, efficient, and privacy preserved. Its prediction accuracy is much better than the models trained from limited personal data and is similar to, and even slightly better than, the results from a centralized dataset. This work paves the way for international collaborations on healthcare projects, where additional data is crucial for reducing bias and providing benefits to humanity.
@@ I Introduction
To address the prevalent challenge of data privacy and access, we use a privacy-preserving framework for AI modeling that leverages federated learning and blockchain. We selected blood glucose prediction as a case study to illustrate the potential of this approach in healthcare applications. Approximately 422 million people globally suffer from diabetes, with 1.5 million deaths directly attributed to the disease annually [7]. An effective strategy for mitigating severe complications due to hyperglycemia or hypoglycemia involves early intervention to regulate blood glucose levels for individuals with type 1 diabetes (T1D). Usually the data obtained from clinical trials are typically scarce and costly to acquire [8]. Research indicates that universal models, trained on data pooled from groups of T1D patients, can significantly enhance glucose prediction accuracy [9].
@@ II-C Blockchain-Enabled FL
Blockchain-enabled FL: Blockchain-enabled FL (BCFL) represents an innovative paradigm of blockchain technology with FL, aiming to enhance security and trust in decentralized machine learning environments. In this paradigm, blockchain serves as an immutable ledger, recording transactions and models exchanged across the distributed nodes participating in the federated learning process. This integration addresses core challenges such as data privacy, security, and model integrity, by ensuring transparent and verifiable transactions while maintaining the confidentiality of the data. Blockchain’s decentralized nature allows for a trustless system where no single entity has control over the entire dataset or the learning process, thereby mitigating risks associated with centralized data storage and processing. Moreover, the use of smart contracts automates the process of data sharing and model updates in FL. Additionally, blockchain’s inherent incentive mechanisms reward honest participation and penalize malicious activities, enhancing further the security of FL systems. BCFL has garnered significant attention for its applicability in diverse fields, including the Internet of Things (IoT) [13, 14] and healthcare [15, 16, 17].
@@ IV-A Experiment Setup
In each experiment, five hospitals participate in a federated learning (FL) training task. Each hospital maintains four weeks of glucose-related data for their Type 1 Diabetes (T1D) patients. Data from one week is used for training, with 5% of this training data set aside for validation. The remaining three weeks of data are used for testing. Overall, we have data from 30 patients, with five patients assigned to each hospital (patient IDs 1 to 25), and five additional patients used exclusively for testing (patient IDs 26 to 30). Specifically, we employed seven methods to train the model: H1Single to H5Single, TotalCentral, and MCGP. H1Single represents single-node training by hospital 1 using its own five patients’ data, TotalCentral involves centralized training using data from 25 patients across all five hospitals, and MCGP refers to blockchain-based federated learning where five hospitals collaboratively train a model using data from these 25 patients. We set the number of previous time steps to 24 and the prediction window to 6, representing a total of 30 minutes. Additionally, in the experiment noted with w/ mal, we included an extra hospital acting as a malicious participant with data from five fake patients.
We utilize LSTM [20] and NNPG [21] models for our experiments to enable edge computing on local hospital or wearable devices. This light configuration not only enhances performance but also significantly reduces the training load, making the process more efficient and feasible for widespread use in patient care. The learning rate for all experiments is set to 0.000001, using the Adam optimizer with a weight decay of 0.0004. In both centralized and federated training, the local epoch is defined as 5000, making sure the local model can be updated to convergences. For FL, we conduct 40 communication rounds to ensure the global model can fully learn the data features from all clients. In each round, 50% of the total participants are assigned the role of proposer, while the remaining participants are assigned the role of voter. Additionally, we use Chainlink oracles  [^2] https://docs.chain.link/vrf to provide the Verifiable Random Function (VRF) functionality.
@@ IV-C Evaluation Metrics
We use root mean square error (RMSE) and mean absolute relative difference (MARD) for BGLP, as they are most widely used metrics to measure glucose accuracy.
@@
where $\hat{g}_{g|k-PH)}$ denotes the prediction results provided the historical data and $y$ denotes the reference glucose measurement, and $N$ refers to the data size.
@@ IV-D1 MCGP Achieves Advanced Model Performance by Breaking Data Silos
To demonstrate the performance of MCGP in breaking data silos by integrating the federated learning paradigm, we compared it with standalone single-node training (SNT) conducted by five other hospitals, labeled H1Single to H5Single. The RMSE and MARD losses for selected current patients are reported in Tables I and II. Our results consistently show that MCGP outperforms the other SNT methods, and even centralised training (TotalCentral) achieving the lowest RMSE and MARD losses both individually and on average. Besides, MCGP’s ability to aggregate global features from all participants enables it to predict more accurately than models trained solely on local data. As illustrated in Figure 4, Participant23’s data, which is part of the H5Single training dataset, does not perform as well as MCGP, particularly at the inflection points of the curve.

### v3

@@ Abstract
One of the biggest challenges of building artificial intelligence (AI) model in the healthcare area is the data sharing. Since healthcare data is private, sensitive, and heterogeneous, collecting sufficient data for modelling is exhausting, costly, and sometimes impossible. In this paper, we propose a framework for global healthcare modelling using datasets from multi-continents (Europe, North America, and Asia) without sharing the local datasets, and choose glucose management as a study model to verify its effectiveness. Technically, blockchain-enabled federated learning is implemented with adaptation to meet the privacy and safety requirements of healthcare data, meanwhile, it rewards honest participation and penalizes malicious activities using its on-chain incentive mechanism. Experimental results show that the proposed framework is effective, efficient, and privacy-preserving. Its prediction accuracy consistently outperforms models trained on limited personal data and achieves comparable or even slightly better results than centralized training in certain scenarios, all while preserving data privacy. This work paves the way for international collaborations on healthcare projects, where additional data is crucial for reducing bias and providing benefits to humanity.
@@ I Introduction
To address the prevalent challenge of data privacy and access, we use a privacy-preserving framework for AI modeling that leverages federated learning and blockchain. We selected blood glucose prediction as a case study to illustrate the potential of this approach in healthcare applications. Approximately 422 million people globally suffer from diabetes, with 1.5 million deaths directly attributed to the disease annually [7]. An effective strategy for mitigating severe complications due to hyperglycemia or hypoglycemia involves early intervention to regulate blood glucose levels for individuals with type 1 diabetes (T1D). Usually, the data obtained from clinical trials are typically scarce and costly to acquire [8]. Research indicates that universal models, trained on data pooled from groups of T1D patients, can significantly enhance glucose prediction accuracy [9].
@@ II-C Blockchain-Enabled FL
Blockchain-enabled FL: Blockchain-enabled FL (BCFL) represents an innovative paradigm of blockchain technology with FL, aiming to enhance security and trust in decentralized machine learning environments. In this paradigm, blockchain serves as an immutable ledger, recording transactions and models exchanged across the distributed nodes participating in the federated learning process. This integration addresses core challenges such as data privacy, security, and model integrity, by ensuring transparent and verifiable transactions while maintaining the confidentiality of the data. Blockchain’s decentralized nature allows for a trustless system where no single entity has control over the entire dataset or the learning process, thereby mitigating risks associated with centralized data storage and processing. Moreover, the use of smart contracts automates the process of data sharing and model updates in FL. Additionally, blockchain’s inherent incentive mechanisms reward honest participation and penalize malicious activities, further enhancing the security of FL systems. BCFL has garnered significant attention for its applicability in diverse fields, including the Internet of Things (IoT) [13, 14] and healthcare [15, 16, 17].
@@ IV-A Experiment Setup
In each experiment, five hospitals participate in a federated learning (FL) training task. Each hospital maintains four weeks of glucose-related data for its Type 1 Diabetes (T1D) patients. Data from one week is used for training, with 5% of this training data set aside for validation. The remaining three weeks of data are used for testing. Overall, we have data from 30 patients, with five patients assigned to each hospital (patient IDs 1 to 25), and five additional patients used exclusively for testing (patient IDs 26 to 30). Specifically, we employed seven methods to train the model: H1Single to H5Single, TotalCentral, and MCGP. H1Single represents single-node training by hospital 1 using its own five patients’ data, TotalCentral involves centralized training using data from 25 patients across all five hospitals, and MCGP refers to blockchain-based federated learning where five hospitals collaboratively train a model using data from these 25 patients. We set the number of previous time steps to 24 and the prediction window to 6, representing a total of 30 minutes. Additionally, in the experiment noted with w/ mal, we included an extra hospital acting as a malicious participant with data from five fake patients.
We utilize LSTM [20] and NNPG [21] models for our experiments to enable edge computing on local hospital or wearable devices. This light configuration not only enhances performance but also significantly reduces the training load, making the process more efficient and feasible for widespread use in patient care. The learning rate for all experiments is set to 0.000001, using the Adam optimizer with a weight decay of 0.0004. In both centralized and federated training, the local epoch is defined as 5000, making sure the local model can be updated to converge. For FL, we conduct 40 communication rounds to ensure the global model can fully learn the data features from all clients. In each round, 50% of the total participants are assigned the role of proposer, while the remaining participants are assigned the role of voter. Additionally, we use Chainlink oracles  [^2] https://docs.chain.link/vrf to provide the Verifiable Random Function (VRF) functionality.
@@ IV-C Evaluation Metrics
We use root mean square error (RMSE) and mean absolute relative difference (MARD) for BGLP, as they are the most widely used metrics to measure glucose accuracy.
@@
where $\hat{g}_{g|k-PH)}$ denotes the prediction results provided by the historical data, and $y$ denotes the reference glucose measurement, and $N$ refers to the data size.
@@ IV-D1 MCGP Achieves Advanced Model Performance by Breaking Data Silos
To demonstrate the performance of MCGP in breaking data silos by integrating the federated learning paradigm, we compared it with standalone single-node training (SNT) conducted by five other hospitals, labeled H1Single to H5Single. The RMSE and MARD losses for selected current patients are reported in Tables I and II. Our results consistently show that MCGP outperforms the other SNT methods, and even centralised training (TotalCentral), achieving the lowest RMSE and MARD losses both individually and on average. Besides, MCGP’s ability to aggregate global features from all participants enables it to predict more accurately than models trained solely on local data. As illustrated in Figure 4, Participant23’s data, which is part of the H5Single training dataset, does not perform as well as MCGP, particularly at the inflection points of the curve.